The probability density function of X is given by f(x)={ 6x(1-x) 0<x<1 0<y<1

{0 otherwise

find the pdf of Y=X3

1
Expert's answer
2021-10-17T18:04:20-0400

"f(x)=6x(1-x),\space 0\lt x\lt1"

"0 , \space elsewhere"

And we have to find the "pdf" "Y=X^3"

Let "G(y)" be the value of the distribution function of "Y" at point "y", it is defined by,

"G(y)=p(Y\leqslant y)"

"=p(X^3\leqslant y)"

"=p(X\leqslant y^{1/3})"

"=\int^{y^{1/3}}_06x(1-x)dx"


"=\int^{y^{1/3}}_0(6x-6x^2)dx"

"=3x^2-2x^3|^{y^{1\/3}}_0"

"=3y^{2\/3}-2y"

In order to find the pdf of random variable "Y" given by "g(y)", then we differentiate "G(y)" with respect to "y" as below.

"g(y)=G'(y)=dG(y)\/dy"

"=d/dy\space (3y^{2/3}-2y)"

"=2y^{-1\/3}-2=2(y^{-1\/3}-1)"

Therefore, the "pdf" of "Y=X^3" is given as,

"g(y)=2(y^{-1/3}-1),\space 0\lt y\lt 1"

"0,\space elsewhere"